# Scalable Training of Graph Foundation Models for Atomistic Materials Modeling: A Case Study with HydraGNN

### v1

@@ Abstract
We present our work on developing and training scalable graph foundation models (GFM) using HydraGNN, a multi-headed graph convolutional neural network architecture. HydraGNN expands the boundaries of graph neural network computations in both training scale and data diversity. It abstracts over message passing algorithms, allowing both reproduction of and comparison across algorithmic innovations that define nearest-neighbor convolution in graph neural networks. This work discusses a series of optimizations that have allowed scaling up the GFM training to tens of thousands of GPUs on datasets that consist of hundreds of millions of graphs. Our GFMs use multi-task learning (MTL) to simultaneously learn graph-level and node-level properties of atomistic structures, such as the total energy and atomic forces. Using over 150 million atomistic structures for training, we illustrate the performance of our approach along with the lessons learned on two state-of-the-art United States Department of Energy (US-DOE) supercomputers, namely the Perlmutter petascale system at the National Energy Research Scientific Computing Center and the Frontier exascale system at Oak Ridge National Laboratory. The HydraGNN architecture enables the GFM to achieve near-linear strong scaling performance using more than 2,000 GPUs on Perlmutter and 16,000 GPUs on Frontier. Hyperparameter optimization (HPO) was performed on over 64,000 GPUs on Frontier to select GFM architectures with high accuracy. Early stopping was applied on each GFM architecture for energy awareness in performing such an extreme-scale task. The training of an ensemble of highest-ranked GFM architectures continued until convergence to establish uncertainty quantification (UQ) capabilities with ensemble learning. Our contribution opens the door for rapidly developing, training, and deploying GFMs using large-scale computational resources to enable AI-accelerated materials discovery and design.
@@ I Introduction
However, generating vast volumes of experimental and/or first-principles data is impractical even with sophisticated experimental facilities and powerful supercomputers. Recently, foundation models (FMs) have demonstrated the capability to navigate around the challenges: once pre-trained over a large volume of available open-source data [15], an FM holds the promise to overcome this limitation by providing a jump-start to refined models by fine-tuning on smaller amounts of data for customized applications (also called downstream tasks). Reducing the number of simulations and/or experiments for generating domain-specific training data also drastically reduces the energy costs of developing domain-specific DL models.
While state-of-the-art language-based FMs with a transformer architecture have reached outstanding results in several domains [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31], they are still ineffective for atomistic materials modeling because they do not capture important topological aspects of the atomistic structures. Therefore, alternative DL architectures that are better suited to retain important topological aspects at the atomistic scale need to be be considered for the development of trustworthy FMs for materials using atomistic scale information.
@@
In this work, we describe our approach towards developing trustworthy supervised GFMs for atomistic materials modeling for simultaneous prediction of energies and atomic forces. Trustworthiness is obtained (i) by performing HPO at extreme scale to identify a sufficiently broad set of architectures capable of reaching the desired accuracy and (ii) by fully training these HPO candidates to obtain high confidence over the predictions using ensemble UQ. The GFMs have been constructed using HydraGNN [36], a fully scalable ORNL-branded architecture. In addition to full scalability [37], HydraGNN offers several other important capabilities: (i) multi-task learning (MTL) for simultaneous prediction of multiple properties and stabilization of the training [38], which we leverage for simultaneous prediction of energies and atomic forces; (ii) object-oriented programming for message passing neural network (MPNN) layers that allows for automated search of the best performing MPNN by treating the choice of the MPNN as a hyperparameter [39]; (iii) invariant and equivariant features that reduce the computational redundancy and time-to-solution [40], therefore saving energy; (iv) scalable input/output (I/O) data management for efficient DDP on supercomputing facilities [41]. In addition, for this work we have added other important capabilities to HydraGNN, such as (a) integration of scalable hyperparameter optimization (HPO) to identify the best-performing configurations of hyperparameters in a computationally efficient manner and (b) scalable ensemble UQ to assess the confidence level of the GFM predictions.
@@
The rest of the paper is organized as follows. We discuss the current state of the art and introduce the HydraGNN architecture in Section II. In Section III, we detail our contributions towards obtaining a scalable framework and list the different optimizations performed for scalable training. We discuss our use of large-scale HPO to develop a trustworthy GFM. Section IV shows the performance of different components of this work: reading large data, scaling the training process, and performing HPO at large scale. We conclude our study and discuss future work in Section V.
@@ III Our Contribution
The work described in this manuscript is the first-of-its-kind large-scale training of GFMs for atomistic materials modeling. Three key innovations of our work are: 1) scalable data management using a scientific data management library and an in-memory data store, 2) scalable HPO that uses asynchronous Bayesian optimization for efficiently managing computing resources, and 3) the use of ensemble methods for uncertainty quantification, which improves model generalization and enables concurrently training multiple models. These three advancements collectively enhance the robustness, efficiency, and scalability of the GNN training process.
@@ III-A Data Aggregation
Each dataset is unique for the chemical compositions and the number of atoms in the atomistic structures of the compounds described. Fig. 1 shows the distribution of the number of atoms and bonds per molecule for each dataset. For the MPTrj dataset, approximately half of the small molecules are relatively small in size. On the other hand, the OC2020 and OC2022 datasets consist of a more even distribution of molecules with different sizes and edge counts, with the larger molecules consisting of over 400 atoms and over 12,500 edges. In total, the data used for training our GFM consisted of 155 million molecules that consume 5.3 Terabytes of storage space. As discussed in the next section, these datasets were pre-processed using a scientific data management library into a common format for efficient storage and I/O.
@@
III-B Scalable Data Management
HydraGNN implements two optimization strategies that address scalability issues due to the large volume of data used for training. These strategies target 1) efficient storage and performant reading of large training data and 2) fast reading of batch data during the training process. As molecular datasets are typically exported as collections of large numbers of files, storing them on the file system and then reading data from the file system during the training process causes a severe I/O bottleneck for GNN training. A large number of small files (e.g., thousands of thousands) puts significant pressure on the metadata service, further slowing data access. Additionally, frequent data fetching by multiple GPUs from the file system during training loops results in a substantial slowdown in the training process. We adopted a two-pronged approach to managing large data and reducing the I/O overhead for training the GNN model. First, we pre-process the various input datasets and store their graph representation using a scientific data management library. Secondly, we use a distributed in-memory data store to load data into memory for fast shuffling of data objects during the training process.
III-B1 ADIOS for High Performance I/O
@@ III-B2 DDStore
DDStore provides options to tune the size of data chunks stored on each process (chunking), replicating a dataset on internal sub-groups of processes (replication), and the communication mechanism selected for fetching data. For our experiments, data is split evenly amongst all processes, and a single replica of the dataset is maintained across all processes. For efficient data retrieval, the low latency MPI one-sided remote memory access (RMA) operations were used. Figure 2 shows the data loading and caching approach used by DDStore compared to traditional approaches that read data directly from the file system. Section IV shows the time taken to obtain a batch of data samples for different model sizes and node counts.
III-C Scalable HPO
@@
III-D Scalable Uncertainty Quantification with GNN Ensembles
@@ IV-A Setup
We utilize three different sizes of foundation models for scaling measurements, denoted as SMALL, MEDIUM, and LARGE. They differ in the total number of parameters, ranging from approximately 60,000 to 163 million. Table II provides details about the three model sizes.
@@
Performance Analysis Tools
Rocprof
AMD’s rocprof is a tool for profiling ML and HPC applications on AMD ROCm platforms such as ORNL’s Frontier. The rocprof tool can be used to identify performance bottlenecks in applications and to optimize their performance. It provides a variety of profiling data, including performance counters, hardware traces, and runtime API/activity traces with start/end timestamps for each API call and GPU kernel execution. The rocprof tool is included as a standard component of the ROCm distribution. The rocprof tool allows the user to choose the specific performance counters to monitor and consolidates all the collected metrics across MPI ranks into a single CSV file.
Real-time GPU Telemetry
To aid in monitoring HydraGNN execution in real-time for each job on Frontier, an AMD research utility was also used to sample a variety of GPU telemetry metrics including occupancy, high-bandwidth memory (HBM) usage, power, temperature, and frequency on a per GCD basis across all nodes assigned to an individual run. This Python-based utility was executed entirely in user-space implemented as a prometheus client on each assigned compute node and combines low-overhead sampling via AMD’s system management interface (SMI) at fixed intervals with a temporary prometheus server [85] instantiated on one CPU core of the master compute host per batch job. Minimal job overhead (less than 0.5%) was observed when running HydraGNN training with this approach for sampling intervals down to 1 sec.
@@
TABLE II: GNN model sizes used for strong and weak scaling tests on NERSC-Perlmutter and OLCF-Frontier.
@@ IV-B I/O Performance for Reading Large Data
In Section III-B, we described using the ADIOS scientific data management library for fast storage and retrieval of large training data. In this section, we show the performance of reading large data in HydraGNN for training models.
@@
Figure 3 shows the reading performance of trainset data on Frontier when all processes read their assigned graphs in parallel. We obtain over 8 Terabytes/second for higher node counts and almost 2 Terabytes/second on 128 nodes. The high I/O bandwidth is a characteristic feature of the ADIOS library as it permits multiple processes to read data spread over multiple ADIOS sub-files efficiently. We omit results from Perlmutter for this part due to filesystem issues encountered on the system during our study.
@@
The initial step in which the root process reads several small portions of the dataset and broadcasts them is an inherently sequential set of operations. As this slows the overall I/O, we obtain lower I/O bandwidth as the root process performs these tasks for the trainset, valset, and testset data groups to read a total of approximately 500 Gigabytes of initial data. Figure 4 shows the sustained I/O bandwidth achieved when HydraGNN reads the entire OC2020 dataset, which is 4.3 Terabytes in size. We obtain a net bandwidth of over 120 Gigabytes/second on Frontier, which allows HydraGNN to ingest the full collection of 120 million graphs in just over 30 seconds.
@@ IV-C HydraGNN Training Scaling Results
We now analyze the scaling performance of HydraGNN on Frontier and Perlmutter. We present weak and strong scaling trends, along with a breakdown of component operations in HydraGNN as we scale it up. Experiments were performed with up to 2,048 nodes on Frontier and 256 nodes on Perlmutter using the three model sizes discussed in Table II.
@@ IV-C1 Weak Scaling
For the weak scaling runs, we configured each GPU to process 3,500 molecules equally. Figure 5 shows the weak scaling performance on Perlmutter and Frontier as we vary in the number of GPUs used for the training. The reported time represents the average training time per epoch. We conducted experiments with up to 2,048 GPUs on Frontier and 1,024 GPUs on Perlmutter. The limited number of GPUs on Perlmutter was due to constraints on available node hours. We observe that the parallel efficiency of weak scaling experiments drops as we increase the number of GPUs beyond 256 for both Perlmutter and Frontier. This is attributed to increased communication costs as we scale the number of GPUs and the overhead associated with using varying graph sizes.
@@
Figure  6 provides a breakdown of the overhead of different components of HydraGNN used in the weak scaling experiments. ‘forward’ and ‘backward’ represent the forward and backward phases of the DL model training, respectively, and ‘dataload’ denotes the cost of obtaining the next batch of data samples from DDStore after a GPU finishes processing its current batch. We notice that ‘dataload’ has a fixed cost, which expectedly becomes more prominent for the small model size and is only a fraction of the runtime as the model size increases. The forward and backward phases show an increase in runtime as we scale up the workflow as synchronization and communication operations become more expensive with increasing GPU counts.
@@ IV-C2 Strong Scaling
For strong scaling runs, we trained HydraGNN on 120 million molecular graphs (approximately 4TB in size) on Frontier and 2 million molecular graphs on Perlmutter for the three model sizes. Figure 7 shows the scaling results for 512 to 16,384 GPUs on Frontier and from 64 up to 2,048 GPUs on Perlmutter. The reported time is the average training time per epoch, similar to the weak scaling measurements. While the SMALL model’s performance deviates from the optimal linear dotted line after 2,048 GPUs on Frontier, the MEDIUM and LARGE models maintain close to linear scaling up to 16,384 GPUs on Frontier. We notice a similar trend on Perlmutter where we observe near-linear scaling upto 2048 GPUs for all model sizes.
@@
The drop in scaling performance is attributed to load imbalance - an artifact of varying graph sizes. As shown in Figure 1, we use a diverse dataset where molecule sizes vary by up to 400, and the number of edges in the larger molecules exceeds 12,500. This results in an imbalanced workload among GPUs in each batch, causing some GPUs to finish training before others. As GPUs must synchronize for exchanging model weights, the runtime is dominated by the GPUs that must train on larger graphs. Effectively, this leads to sub-standard utilization of compute resources and posts a challenge towards achieving high-performant, scalable training. Training on large volumes of data can help develop robust models because of the diverse nature of data, but it affects the computational performance as the workload can vary greatly.
@@
The EGNN model we used is particularly vulnerable to this problem. The time required for forward calculations in EGNN is directly proportional to the number of edges in the molecules. For datasets with highly variable edge counts between molecules, the likelihood of load imbalance between GPUs increases. Figure 8 illustrates the time spent on the forward task in the EGNN model with different model sizes. We observe an almost linear relationship between forward time and graph size at each batch (measured by the number of edges). The SMALL models show large variances on both machines, which is expected due to system noise being more apparent with smaller model sizes. Significant performance differences (e.g., the difference between minimum and maximum time) are observed due to the varying graph sizes in our datasets. However, for other tasks (data loading and backward), we do not observe a similar correlation, as they are agnostic of the graph size. Fig. 9 illustrates the average percentage of time spent waiting during three tasks: data loading, forward pass, and backward pass. It highlights a significant waiting period during the forward pass, primarily due to varying graph sizes. This waiting time increases as the disparity in graph sizes among GPUs grows. Other tasks, such as data loading and backward pass, also involve waiting time, but to a much lesser extent.

### v2

@@ Abstract
We present our work on developing and training scalable graph foundation models (GFM) using HydraGNN, a multi-headed graph convolutional neural network architecture. HydraGNN expands the boundaries of graph neural network computations in both training scale and data diversity. It abstracts over message passing algorithms, allowing both reproduction of and comparison across algorithmic innovations that define nearest-neighbor convolution in graph neural networks. This work discusses a series of optimizations that have allowed scaling up the GFM training to tens of thousands of GPUs on datasets that consist of hundreds of millions of graphs. Our GFMs use multi-task learning (MTL) to simultaneously learn graph-level and node-level properties of atomistic structures, such as the total energy and atomic forces. Using over 150 million atomistic structures for training, we illustrate the performance of our approach along with the lessons learned on two state-of-the-art United States Department of Energy (US-DOE) supercomputers, namely the Perlmutter petascale system at the National Energy Research Scientific Computing Center and the Frontier exascale system at Oak Ridge National Laboratory. The HydraGNN architecture enables the GFM to achieve near-linear strong scaling performance using more than 2,000 GPUs on Perlmutter and 16,000 GPUs on Frontier. Hyperparameter optimization (HPO) was performed on over 64,000 GPUs on Frontier to select GFM architectures with high accuracy. Early stopping was applied on each GFM architecture for energy awareness in performing such an extreme-scale task. The training of an ensemble of highest-ranked GFM architectures continued until convergence to establish uncertainty quantification (UQ) capabilities with ensemble learning. Our contribution establishes core capabilities for rapidly developing, training, and deploying further GFMs using large-scale computational resources to enable AI-accelerated materials discovery and design.
@@ I Introduction
However, generating vast volumes of experimental and/or first-principles data is impractical even with sophisticated experimental facilities and powerful supercomputers. Recently, foundation models (FMs) have demonstrated the capability to navigate around the challenge: once pre-trained over a large volume of available open-source data [15], an FM holds the promise to overcome this limitation by providing a jump-start to refined models by fine-tuning on smaller amounts of data for customized applications (also called downstream tasks). Reducing the number of simulations and/or experiments for generating domain-specific training data also drastically reduces the energy costs of developing domain-specific DL models.
While state-of-the-art language-based FMs with a transformer architecture have reached outstanding results in several domains [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31], they fail to capture important topological aspects of the atomistic structures. Therefore, alternative DL architectures that are better suited to retain important topological aspects at the atomistic scale need to be be considered for the development of trustworthy FMs for materials using atomistic scale information.
@@
The work described in this manuscript is the first-of-its- kind large-scale training of GFMs for atomistic materials modeling. We describe our approach to developing trustworthy supervised GFMs for atomistic materials modeling for the simultaneous prediction of energies and atomic forces. Trustworthiness is obtained (i) by performing HPO at extreme scale to identify a sufficiently broad set of architectures capable of reaching the desired accuracy and (ii) by fully training these HPO candidates to obtain high confidence over the predictions using ensemble UQ. The GFMs have been constructed using HydraGNN [36], a fully scalable GNN architecture developed at ORNL. In addition to full scalability [37], HydraGNN offers several other important capabilities: (i) multi-task learning (MTL) for simultaneously predicting multiple properties and stabilizing the training [38], which we leverage for the simultaneous prediction of energies and atomic forces; (ii) an object-oriented design for message passing neural network (MPNN) layers that allows for automated search of the best performing MPNN by treating the choice of the MPNN as a hyperparameter [39]; (iii) invariant and equivariant features that reduce the computational redundancy and time-to-solution [40], therefore saving energy; and (iv) scalable input/output (I/O) data management for efficient DDP on supercomputing facilities [41]. In addition, for this work we have added important capabilities to HydraGNN for the (a) integration of scalable hyperparameter optimization (HPO) to identify the best-performing configurations of hyperparameters in a computationally efficient manner and (b) scalable ensemble UQ to assess the confidence level of the GFM predictions.
@@
The rest of the paper is organized as follows. We discuss the current state of the art and introduce HydraGNN in Section II. In Section III, we discuss our approach toward developing a scalable framework and list the different optimizations for scalable training. We discuss our use of large-scale HPO to develop a trustworthy GFM. Section IV shows the performance of different components of this work: reading large data, scaling the training process, and performing HPO at large scale. We conclude our study and discuss future work in Section V.
@@ III Our Contribution
The work described in this manuscript is the first-of-its-kind large-scale training of GFMs for atomistic materials modeling. We have employed three key techniques for developing a scalable and trustworthy GFM: 1) scalable data management using a scientific data management library and an in-memory data store, 2) scalable HPO that uses asynchronous Bayesian optimization for efficiently managing computing resources, and 3) ensemble methods for uncertainty quantification that allows model generalization and concurrently training multiple models. These three advancements collectively enhance the robustness, efficiency, and scalability of the GNN training process.
@@ III-A Data Aggregation
Each dataset is unique for the chemical compositions and the number of atoms in the atomistic structures of the compounds described. Fig. 1 shows the distribution of the number of atoms and bonds per molecule for each dataset. For the MPTrj dataset, approximately half of the molecules are relatively small in size. On the other hand, the OC2020 and OC2022 datasets consist of a more even distribution of molecules with different sizes and edge counts, with the larger molecules consisting of over 400 atoms and over 12,500 edges. In total, the data used for training our GFM consisted of 155 million molecules that consume 5.3 Terabytes of storage space. These datasets were pre-processed using a scientific data management library into a common format for efficient storage and I/O, as discussed in Section III-C.
@@
III-B Data Cleaning
Some of the atomistic structures were determined to have unrealistic values for atomic forces, in the order of 20,000 eV/angstrom. Thus, we first applied a data cleaning task in which we discarded all atomistic structures with an $L2$-norm (also known as spectral norm) of the force tensor above 100 eV/angstrom to ensure that these data samples did not affect the training of our GFMs. The number of data samples removed from each dataset by this filtering operation is reported in Table II.
Dataset Number of data samples removed
ANI1x [65] 0
QM7-X [66] 0
OC2020 [44] 1
OC2022 [45] 12,270
MPTrj [43] 151
Total 12,421
TABLE II: Number of data samples discarded from each dataset in which the $L2$-norm (also known as spectral norm) of the force tensor was over 100 eV/angstrom.
III-C Scalable Data Management
HydraGNN implements two optimization strategies that address scalability issues due to the large volume of data used for training. These strategies aim for 1) efficient storage and performant reading of large training data, and 2) fast reading of batch data during the training process. As molecular datasets are typically exported as collections of large numbers of files, storing datasets on a shared file system and then reading data from the large number of files during the training process causes a severe I/O bottleneck for GNN training. Multiple datasets cumulatively containing tens of thousands of small files put significant pressure on the filesystem’s metadata service, further slowing data access. Additionally, frequent data fetching by multiple GPUs from the file system during training loops results in a substantial slowdown in the training process. We adopted a two-pronged approach to managing large data and reducing the I/O overhead for training the GNN model. First, we pre-process the various input datasets and store their graph representation using a scientific data management library. Secondly, we use a distributed in-memory data store to load data into memory for fast shuffling of data objects during the training process.
III-C1 ADIOS for High Performance I/O
@@ III-C2 DDStore
DDStore provides options to tune the size of data chunks stored on each process (chunking), replicating a dataset on internal sub-groups of processes (replication), and the communication mechanism selected for fetching data. For our experiments, data is split evenly amongst all processes, and a single replica of the dataset is maintained across all processes. For efficient data retrieval, the low latency MPI one-sided remote memory access (RMA) operations were used. Fig. 2 shows the data loading and caching approach used by DDStore compared to traditional approaches that read data directly from the file system. Section IV shows the time taken to obtain a batch of data samples for different model sizes and node counts.
III-D Scalable HPO
@@
III-E Scalable Uncertainty Quantification with GNN Ensembles
@@ IV-A Setup
We utilize three different sizes of foundation models for scaling measurements, denoted as SMALL, MEDIUM, and LARGE. They differ in the total number of parameters, ranging from approximately 60,000 to 163 million. Table III provides details about the three model sizes.
@@
To aid in monitoring HydraGNN execution in real-time for a subset of the analysis carried out on Frontier, an AMD research utility, omnistat, was employed to sample a variety of GPU telemetry metrics including occupancy, high-bandwidth memory (HBM) usage, power, temperature, and frequency on a per GCD basis across all nodes assigned to an individual run. This Python-based utility was executed entirely in user-space implemented as a prometheus client on each assigned compute node and combines low-overhead sampling via AMD’s system management interface (SMI) at fixed intervals with a temporary prometheus server [85] instantiated on one CPU core of the master compute host per batch job. Minimal job overhead (less than 0.5%) was observed when running HydraGNN training with this approach for sampling intervals down to one second.
@@
TABLE III: GNN model sizes used for strong and weak scaling tests on NERSC-Perlmutter and OLCF-Frontier.
@@ IV-B I/O Performance for Reading Large Data
In Section III-C, we described using the ADIOS scientific data management library for fast storage and retrieval of large training data. In this section, we show the performance of reading large data in HydraGNN for training models.
@@
Fig. 3 shows the reading performance of trainset data on Frontier when all processes read their assigned graphs in parallel. We obtain over 8 Terabytes/second for higher node counts and almost 2 Terabytes/second on 128 nodes. The high I/O bandwidth is a characteristic feature of the ADIOS library as it permits multiple processes to read data spread over multiple ADIOS sub-files efficiently. A similar run on Perlmutter was not possible due to filesystem issues encountered on the system during our study.
@@
The initial step in which the root process reads several small portions of the dataset and broadcasts them is an inherently sequential set of operations. As this slows the overall I/O, we obtain lower I/O bandwidth as the root process performs these tasks for the trainset, valset, and testset data groups to read a total of approximately 500 Gigabytes of initial data. Fig. 4 shows the sustained I/O bandwidth achieved when HydraGNN reads the entire OC2020 dataset, which is 4.3 Terabytes in size. We obtain a net bandwidth of over 120 Gigabytes/second on Frontier, which allows HydraGNN to ingest the full collection of 120 million graphs in just over 30 seconds.
@@ IV-C HydraGNN Training Scaling Results
We now analyze the scaling performance of HydraGNN on Frontier and Perlmutter. We present weak and strong scaling trends, along with a breakdown of component operations in HydraGNN as we scale it up. Experiments were performed with up to 2,048 nodes on Frontier and 256 nodes on Perlmutter using the three model sizes discussed in Table III.
@@ IV-C1 Weak Scaling
For the weak scaling runs, we configured each GPU to process 3,500 molecules equally. Fig. 5 shows the weak scaling performance on Perlmutter and Frontier as we vary in the number of GPUs used for the training. The reported time represents the average training time per epoch. We conducted experiments with up to 2,048 GPUs on Frontier and 1,024 GPUs on Perlmutter. The limited number of GPUs on Perlmutter was due to constraints on available node hours. We observe that the parallel efficiency of weak scaling experiments drops as we increase the number of GPUs beyond 256 for both Perlmutter and Frontier. This is attributed to increased communication costs as we scale the number of GPUs and the overhead associated with using varying graph sizes.
@@
Fig.  6 provides a breakdown of the overhead of different components of HydraGNN used in the weak scaling experiments. ‘forward’ and ‘backward’ represent the forward and backward phases of the DL model training, respectively, and ‘dataload’ denotes the cost of obtaining the next batch of data samples from DDStore after a GPU finishes processing its current batch. We notice that ‘dataload’ has a fixed cost, which expectedly becomes more prominent for the small model size and is only a fraction of the runtime as the model size increases. The forward and backward phases show an increase in runtime as we scale up the workflow as synchronization and communication operations become more expensive with increasing GPU counts.
@@ IV-C2 Strong Scaling
For strong scaling runs, we trained HydraGNN on 120 million molecular graphs (approximately 4TB in size) on Frontier and 2 million molecular graphs on Perlmutter for the three model sizes. Fig. 7 shows the scaling results for 512 to 16,384 GPUs on Frontier and from 64 up to 2,048 GPUs on Perlmutter. The reported time is the average training time per epoch, similar to the weak scaling measurements. While the SMALL model’s performance deviates from the optimal linear dotted line after 2,048 GPUs on Frontier, the MEDIUM and LARGE models maintain close to linear scaling up to 16,384 GPUs on Frontier. We notice a similar trend on Perlmutter where we observe near-linear scaling upto 2048 GPUs for all model sizes.
@@
The drop in scaling performance is attributed to load imbalance - an artifact of varying graph sizes. As shown in Fig. 1, we use a diverse dataset where molecule sizes vary by up to 400, and the number of edges in the larger molecules exceeds 12,500. This results in an imbalanced workload among GPUs in each batch, causing some GPUs to finish training before others. As GPUs must synchronize for exchanging model weights, the runtime is dominated by the GPUs that must train on larger graphs. Effectively, this leads to sub-standard utilization of compute resources and posts a challenge towards achieving high-performant, scalable training. Training on large volumes of data can help develop robust models because of the diverse nature of data, but it affects the computational performance as the workload can vary greatly.
@@
The EGNN model we used is particularly vulnerable to this problem. The time required for forward calculations in EGNN is directly proportional to the number of edges in the molecules. For datasets with highly variable edge counts between molecules, the likelihood of load imbalance between GPUs increases. Fig. 8 illustrates the time spent on the forward task in the EGNN model with different model sizes. We observe an almost linear relationship between forward time and graph size at each batch (measured by the number of edges). The SMALL models show large variances on both machines, which is expected due to system noise being more apparent with smaller model sizes. Significant performance differences (e.g., the difference between minimum and maximum time) are observed due to the varying graph sizes in our datasets. However, for other tasks (data loading and backward), we do not observe a similar correlation, as they are agnostic of the graph size. Fig. 9 illustrates the average percentage of time spent waiting during three tasks: data loading, forward pass, and backward pass. It highlights a significant waiting period during the forward pass, primarily due to varying graph sizes. This waiting time increases as the disparity in graph sizes among GPUs grows. Other tasks, such as data loading and backward pass, also involve waiting time, but to a much lesser extent.